# thanks: This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [RSF]

### v2

@@ III.2 Emittance Compensation Using a Solenoidal Lens
A solenoidal lens is installed near the cathode (Fig. 4) to counteract early nonlinear space‐charge effects. The on‐axis field profile used in the simulations (Fig. 5) was optimized via iterative scans, yielding a peak of $B_{z}=125$ mT at $z=85$ mm downstream of the photocathode.
@@
Fig. 10(a) shows the resulting RMS transverse beam size. Immediately after the RF cavities, the solenoid collimates the envelope, stabilizing $\sigma_{x,y}$ along the beamline. The normalized emittance (Fig. 10(b)) initially rises due to the non-zero solenoidal field at the cathode, which imparts extrinsic angular momentum onto the beam. In addition, the combination of the solenoidal field with the transverse RF fields of the 1.5-cell $\pi$-mode gun modifies the evolution of the transverse phase space and suppresses the natural cancellation of RF-induced correlations that occurs in the zero-solenoid case [Floettmann15].
@@
Figure 10: Simulated beam dynamics using the emittance compensation solenoid. (a): Simulated RMS transverse beam size along the beamline with the compensation solenoid. The size stabilizes soon after the RF cavities, indicating effective collimation. (b): Simulated normalized transverse emittance along the beamline with solenoid compensation. Stabilizes near $3.2\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$ downstream.
Parametric scans further validate the robustness of the chosen settings: Fig. 11(a) confirms the optimum solenoid peak at z=85z=85\, mm (fixed $B_{z}=125\,$ mT), and Fig. 11(b) shows that $B_{z}=125\,$ mT yields the minimum final emittance when positioned at this location. These settings provide a stable and experimentally accessible working point for the injector under the present operating conditions.
Figure 11: Final normalized transverse emittance versus solenoid parameters: (a) dependence on the solenoid peak position for a fixed field of 125125\,mT; (b) dependence on the peak solenoid field strength with the solenoid fixed at z=85z=85\,$m$. In both cases the minimum emittance is achieved for a peak field of 125125\,mT at z=85z=85\,$m$.
IV Quantum Beam Dynamics
@@ IV.1 Vortex Electrons
One of the mechanisms for generating the vortex electrons that can be realized at an accelerator facility in an RF photoinjector is photoemission driven by structured light. Theory shows that an atom placed on the axis of an optical vortex transfers a well-defined OAM projection to the emitted electron, and this behaviour persists for ensembles [Pavlov24, Kazinski25]. Alternative schemes based on electron diffraction [Uchida10, Verbeeck10, Blackburn14] are impractical in such setups.
IV.2 Quantum Evolution of a Single-Electron Wave Packet in the RF Field
@@
Throughout this section, the angular brackets denote quantum expectation values over the single-electron state. We compare three representative cases::
@@
where for acceleration of electrons along +z+z one has $E<0$ and hence $qE>0$, so that $t(z)>0$. The transverse second moment at a fixed plane $z$ is then obtained by evaluating Eq. (4) at this $t(z)$. For typical DC-gun this results in a $\sim 550$-fold reduction of the final transverse width compared to free-space propagation (Fig. 12), in good agreement with recent relativistic calculations [z6j7-grs2].
@@
In the S-band photoinjector electrons are accelerated in a standing-wave RF field. Along the axis, where the packet propagates, the longitudinal electric field can be written as
@@
For the photoinjector parameters used in this work, the resulting relativistic velocity profile reduces the time of flight to the exit plane by almost three orders of magnitude compared to free-space propagation. In the regime $t\gg t_{d}$, where the transverse width obeys $\sigma_{\perp}(t)\propto t$, this kinematic shortening leads to an $\sim 900$-fold suppression of the transverse rms size compared to the free electron at $z=30~\mathrm{cm}$, as shown in Fig. 13. Importantly, similar suppression of quantum spreading persists for all other parameter sets explored in this work (not shown), indicating that strong reduction of transverse spreading is a robust feature of relativistic acceleration rather than a special property of the specific cases illustrated in Figs. 1312.
@@ V Conclusions
Overall, the simulations show that the present injector configuration satisfies the beam-quality requirements for experiments on the photoemission-driven generation and acceleration of vortex electrons in the MeV energy range. The photoinjector has been assembled, and the first photoemitted electron beam has been produced. Commissioning of the RF gun is ongoing, with work underway to reach the design input power of 6 MW. In parallel, a phase-locking system for synchronizing the UV laser pulses with the RF field is being developed. All preparations for experiments on photoemission-based generation of relativistic vortex electrons have been completed: diffractive optical elements for producing ultraviolet twisted light beams with orbital angular momentum up to $\ell=64\hbar$ have been fabricated and tested. Future studies will extend the present modelling by including collective interactions and decoherence mechanisms to obtain a complete description of structured electron beams under realistic experimental conditions. The check of the electron beams vorticity can be made by diffraction on a triangle apperture, as has been argued in Ref. [Maksimov25].

### v3

@@ 3.3 Emittance compensation using a solenoidal lens
A solenoidal lens is installed near the cathode (see Fig. 5) to counteract early nonlinear space-charge effects. The on-axis field profile used in the simulations (see Fig. 5) was optimized via iterative scans, yielding a peak of $B_{z}=125$ mT at $z=85$ mm downstream of the photocathode.
@@
Fig. 13 shows the resulting RMS transverse beam size. Immediately after the RF cavities, the solenoid collimates the envelope, stabilizing $\sigma_{x,y}$ along the beamline. The normalized emittance (see Fig. 13) initially rises due to the non-zero solenoidal field at the cathode, which imparts extrinsic angular momentum onto the beam. In addition, the combination of the solenoidal field with the transverse RF fields of the 1.5-cell $\pi$-mode gun modifies the evolution of the transverse phase space and suppresses the natural cancellation of RF-induced correlations that occurs in the zero-solenoid case [Floettmann15].
@@
Figure 12: Simulated RMS transverse beam size along the beamline with the compensation solenoid. The size stabilizes soon after the RF cavities, indicating effective collimation.
Figure 13: Simulated normalized transverse emittance along the beamline with solenoid compensation. Stabilizes near $3.2\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$ downstream.
To verify that the cathode solenoid has been properly designed and that its magnetic field provides effective beam stabilization, we performed additional particle-tracking simulations through the RF gun (see Fig. 16). Figure 15 shows the transverse particle distribution at the exit of the accelerating structure. The rms transverse spread does not exceed 5 mm, which allows the beam to be safely transported through the 14-mm-diameter coaxial aperture.
Figure 14: Simulated transverse particle distribution at the exit of the RF accelerating structure in the presence of the cathode solenoid. The transverse spread of the beam remains within 5 mm, allowing transmission through the 14-mm-diameter coaxial aperture.
Figure 15: Simulated transverse phase space of the electron beam at the RF-gun exit. The $xx^{\prime}$ projection exhibits an approximately ideal phase-space ellipse, indicating stable beam formation.
In addition, Fig. 15 shows the two-dimensional transverse phase space of the beam. The $xx^{\prime}$ projection exhibits an almost ideal phase-space ellipse, indicating stable transverse dynamics and good beam quality, even at relatively large transverse emittance values. Since the simulation corresponds to on-axis beam transport through the solenoid, an identical phase-space distribution is expected in the $yy^{\prime}$ plane.
Figure 16 presents the particle trajectories along the beamline. Most particles are transmitted through the RF gun to its exit. A small fraction of electrons remains near the photocathode region, while another fraction is partially lost at the coaxial aperture. Nevertheless, the total beam loss does not exceed 13%, which indicates a satisfactory operating regime of the cathode solenoid and confirms its effectiveness for beam confinement.
Figure 16: Simulated particle trajectories along the RF photogun beamline with the cathode solenoid. Black dots denote transmitted (active) particles, blue dots correspond to particles remaining in the photocathode region, and red dots indicate particles lost inside the RF structure.
4 Quantum beam dynamics
@@ 4.1 Vortex electrons
One mechanism for generating vortex electrons in an RF photoinjector at an accelerator facility is laser-driven emission using structured light. Theory shows that an atom placed on the axis of an optical vortex transfers a well-defined OAM projection to the emitted electron, and this behaviour persists for ensembles [Pavlov24, Kazinski25]. Alternative schemes based on electron diffraction [Uchida10, Verbeeck10, Blackburn14] are impractical in such setups.
4.2 Quantum evolution of a single-electron wave packet in the accelerating field
@@
Throughout this section, the angular brackets denote quantum expectation values over the single-electron state. We compare three representative cases:
@@
where for acceleration of electrons along +z+z one has $E<0$ and hence $qE>0$, so that $t(z)>0$. The transverse second moment at a fixed plane $z$ is then obtained by evaluating Eq. (4) at this $t(z)$. For typical DC-gun this results in a $\sim 550$-fold reduction of the final transverse width compared to free-space propagation (see Fig. 17), in good agreement with recent relativistic calculations [z6j7-grs2].
@@
In the S-band photoinjector electrons are accelerated in a standing-wave RF field. Along the axis, where the packet propagates, the longitudinal electric field can be written as [Wiedemann15, Chao20, Lee21]
@@
For the photoinjector parameters used in this work, the resulting relativistic velocity profile reduces the time of flight to the exit plane by almost three orders of magnitude compared to free-space propagation. In the regime $t\gg t_{d}$, where the transverse width obeys $\sigma_{\perp}(t)\propto t$, this kinematic shortening leads to an $\sim 900$-fold suppression of the transverse rms size compared to the free electron at $z=30~\mathrm{cm}$, as shown in Fig. 18. Importantly, similar suppression of quantum spreading persists for all other parameter sets explored in this work, indicating that strong reduction of transverse spreading is a robust feature of relativistic acceleration rather than a special property of the specific cases illustrated in Figs. 1817.
@@ 5 Conclusions
Overall, the simulations show that the present injector configuration satisfies the beam-quality requirements for experiments on the photoemission-driven generation and acceleration of vortex electrons in the MeV energy range. The photoinjector has been assembled, and the first photoemitted electron beam has been produced. Commissioning of the RF gun is ongoing, with work underway to reach the design input power of 6 MW. In parallel, a phase-locking system for synchronizing the UV laser pulses with the RF field is being developed. All preparations for experiments on photoemission-based generation of relativistic vortex electrons have been completed: diffractive optical elements for producing ultraviolet twisted light beams with orbital angular momentum up to $\ell=64\hbar$ have been fabricated and tested. Future studies will extend the present modelling by including collective interactions and decoherence mechanisms to obtain a complete description of structured electron beams under realistic experimental conditions. The check of the electron beams vorticity can be made by diffraction on a triangle apperture, as has been discussed in Ref. [Maksimov25].